# Localized states, topology and anomalous Hall conductivity on a 30 degrees twisted bilayer honeycomb lattice

### v3

@@ Abstract
We consider $30^{\circ}$ twisted bilayer formed by two copies of Haldane model and explore its evolution with varying interlayer coupling strength. Specifically, we compute the system’s energy spectrum, its fractal dimensions, topological entanglement entropy, local Chern markers and anomalous Hall conductivity. We find that at weak interlayer coupling, the system still has a bulk energy gap, topological edge states and retains topological properties of the isolated layers, but at strong interlayer coupling, this energy gap closes, but it is possible that another bulk gap may open. Such a system possesses multiple states localized at various locations of the lattice. We also establish that topological entanglement entropy and anomalous Hall conductivity can be used to characterize the system’s topological properties in the same way as a local Chern marker.
@@ I Introduction
Motivated by this, we are interested in studying a quasicrystal obtained by smoothly deforming a crystal with known non-trivial topology. Specifically, we choose to consider 30 degrees twisted bilayer honeycomb lattice - a model which is currently actively studied in the context of twisted bilayer graphene [12, 13, 14, 15]. However we consider a twisted bilayer, in which each monolayer is described by gapped Haldane model [16, 17, 18]. We explore our model’s evolution once the interlayer coupling is increased. We obtain that at weak coupling, the model retains all properties of Haldane model: its bulk spectrum remains gapped and the topological edge states persist. Thus we show that our system of two coupled copies of Chern insulators can be viewed as a topological quasicrystal. On the other hand, as the interlayer coupling becomes larger, the bulk energy gap closes, and the topological edge states disappear. Yet, since this system is non-periodic, its eigenstates are generally not extended. We obtain numerically that our bilayer possesses multiple kinds of states localized at various locations of the lattice. These include, but not limited to corner states, which were previously found in Refs. [17, 18]. We also observe that at large interlayer coupling, a new energy gap may open, but we argue that it is not topological.
@@
This paper is organized as follows. In Sec. II we introduce our model, and in Sec. II.1 we describe its phase diagram, which consists of multiple phases: topological weakly-coupled, non-topological weakly coupled and strongly coupled multifractal. The latter may have both gapped and gapless phases. We also comment on multifractal properties of our model (II.2). In Sec. II.3 we describe its topological properties. Specifically, in Sec. II.3.1 we present our results for its topological entanglement entropy and in Sec. II.3.2 we describe its topological properties using the ’local Chern markers’. In Sec. II.3.3 we discuss anomalous Hall conductivity. We summarize our findings in Sec. III .
@@ II Model
$H=\sum\limits_{\alpha}H_{\alpha}+V_{1-2},$ (1)
where $\alpha=1,2$ is the layer index, $H_{\alpha}$ is Haldane Hamiltonian of each layer, and $V_{1-2}$ describes interaction between the layers. Specifically,
@@
Here $i,j$ numerate the lattice sites within each layer, <i,j><i,j> is a sum over the nearest neighbors, $\ll i,j\gg$ is a sum over the next-nearest neighbors. $t_{intra}$ is the nearest neghbors hopping (between A, B sites), $t_{2}$ is the next-nearest neighbors hopping (between A-A or B-B sites), $\nu=\pm 1$ depending on the relative direction between $i,j$, $\phi=\pi/2$, and $\epsilon=\pm 1$ for A(B) sites (see e.g. ref. [24] for more details). The interlayer potential has the form
$v_{i,j}=\left\{e^{-\frac{r_{ij}}{r_{0}}},\quad r<r_{max}\atop 0,\qquad r>r_{max}.\right$. (4)
Here $r_{ij}$ is a distance between the sites $i,j$, and to simplify the model we neglect interlayer spacing. Also we introduced long-distance cutoff $r_{max}$ to decrease the complexity of our numerical calculations. We fix the numerical values of our parameters as $t_{intra}=1$, $t_{2}=1$, $\phi=\pi/2$, $r_{0}=1$, $r_{max}=2$ (we checked that changing $t_{2}$, $r_{max}$ does not have qualitative impact on the model’s properties) and explore its evolution over $m$, $t_{inter}$. Since our 30 degrees twisted bilayer is a non-periodic system, we consider a system forming a square with equal lengths in both directions $L_{x}=L_{y}$ and fix open boundary conditions.
@@
Figure 1: A schematic picture of the $30^{\circ}$ twisted bilayer of two honeycomb lattices. Here we show its top-left corners with $L_{x}=20$ (1(a)), $L_{x}=30$ (1(b)), $L_{x}=40$ (1(c)). The top layer (blue) is formed by sites with coordinates $\vec{r}_{A,B}=\vec{a}_{1}n_{1}+\vec{a}_{2}n_{2}+\vec{c}_{A,B}$, where $\vec{a}_{1}=(\sqrt{3},0)$ , $\vec{a}_{2}=(\sqrt{3}/2,3/2)$, $\vec{c}_{A,B}=(0,\pm 1)$. The bottom layer (red) is obtained by rotating the top layer at 30 degrees. We consider an open system, whose sites are inside a square $L_{x}<x,y<L_{x}$. The cases (1(a), 1(b)) correspond to zigzag-armchair edges, and the case (1(c)) corresponds to the bearded-armchair edge.
@@ II.1 Phase diagram
where $P_{q}$ is a participation ratio defined as $P_{q}=\sum_{i}(\psi^{\dagger}\psi)^{q}$, and $i=1\ldots N$ runs over all lattice sites. In a system of large spatial size, fractal dimension is expected to be close to $1$ for extended bulk states, $1/2$ for edge states and 0 for localized states (e.g. corner modes). In this section, we use $D_{q}$ to find the most localized states, which have the smallest $D_{q}$.
@@
First, let us recall our model’s properties in the known case when the interlayer coupling $t_{inter}$ is set to zero. This model is a Chern insulator at $|m|<1$ (we assume that both layers have the same signs of Chern numbers). Its bulk states form two bands separated by an energy gap, which in turn, is filled by edge states. At $m=1$, the bulk gap closes, and at $m>1$, it reopens again, and the system becomes a trivial insulator.
@@
Once the interlayer coupling is turned on, the phase diagram of the system in $m,t_{inter}$ space has multiple phases, which we schematically outline on the Fig. 2 and which we obtained by running over different values of $m$ and $t_{inter}$ (see Figs. 3, 4). Specifically, if we start from the topological phase of the Haldane model, we observe (see Figs. 3) that at small coupling, the bulk states are still separated by an energy gap and the edge states persist, but as $t_{inter}$ is increasing, the bulk energy gap shrinks. Once $t_{inter}$ reaches a critical value ($t_{inter}\approx 1.5$ for $m=0$), the bulk gap closes, and the edge states disappear. As we increase $t_{inter}$ further, at small $m$, another energy gap appears (see Fig. 4(a), 4(b), 4(c) ). However, we do not observe it at larger $m$ (see Fig. 4(d) ). Similarly, if we start from the non-topological phase (see Fig. 4(f)), at small $t_{inter}$, the system remains gapped, and as $t_{inter}$ is increasing, the gap becomes smaller and eventually disappears. Near this phase transition, another gap may appear.
@@
Let us look more closely at the properties of the states in our model. For simplicity, we focus on the case $m=0$. As we mentioned, at $t_{inter}=0$, our model is a superposition of two Haldane models, which in turn host bulk states separated by an energy gap and edge states within the bulk gap. However, we also found that our system hosts additional edge states within the bulk gap. We observe such states in the cases of both zigzag-armchair and bearded-armchair edges, but we find them to lie at slightly different energies. Specifically, in the case of zigzag-armchair edges, these edge states lie at $E\approx\pm 2.3$ (see Fig. 6(a)), whereas in the case of bearded-armchair edge, these states lie at energy $E\approx\pm 1.9$. One can see these states for various values of $L_{x}$ on the Fig. 5(a) - there bearded-armchair configuration corresponds to lattice sizes $7,10,13,16\ldots$, and zigzag-armchair is realized for all other integer values of $L_{x}$.
@@
As we increase $t_{inter}$, the edge states in the bottom ‘band’ disappear. However, in the top ‘band’, they persist, and a few of them tend to become localized within the corners of our system (see Figs. 6(b), 6(e)). Specifically for $L_{x}=30$, we also find that a few of the topological edge states become localized at the corners (see Fig. 6(b); note that we could not establish connection between these modes and a shape of the lattice boundary). As we increase $t_{inter}$ further, the bulk gap closes, and the system becomes gapless. At even larger $t_{inter}$, another energy gap opens. Interestingly, we found that even though such a gap exists at any system size, its precise location and width varies a lot between different lattice sizes (see Figs. 5(c), 5(d)). We attribute it to the fact that since our system is quasiperiodic, the system becomes unique for each lattice size, and therefore it is extremely difficult to achieve ‘continuum limit’ (see also discussion in the Sec. II.2). We however, can see that as the system size increases, this gap becomes less pronounced (see Fig. 5).
@@
As predicted in the previous works [16], our system hosts states localized at the corners. However, we argue that their location is not attributed to the bulk gap. Indeed, we can see that for $L_{x}=30$ (see Fig. 4(b)), the localized corner states are in the bulk gap for a range of $t_{inter}$ between $\approx 5$ and $10$, but at smaller or larger $t_{inter}$, they merge with the bulk spectrum. On the other hand, at $L_{x}=20$ (see Fig. 4(a)), we find that the corner modes lie entirely within the bulk spectrum away from the gap, but at the same time, within the bulk gap, there are extended states. We note that both $L_{x}=20,30$ in our example correspond to zigzag-armchair edges of the lattice, and even have a similar structure of the corners (see Fig. 1(a), 1(b) ). We observe qualitatively similar picture for $L_{x}=40$ with ‘armchair-bearded’ edge structure (Fig. 1(c)): localized corner modes lie within the bulk gap at $t_{inter}>5$, but within the bulk states at smaller $t_{inter}$. ((see Fig. 4(c))).
@@
Even more interestingly, we observe numerically that our system hosts other kinds of localized states. For instance, even at large $t_{inter}$, our system still hosts states localized near edges (see Fig. 7(f) ). But most surprisingly, we find that, at $m=0$, the states with the smallest fractal dimensions are localized in the center of the lattice (see Fig. 7(a), 7(b)).
@@
We also find that at $m\neq 0$, both corner and ‘center’ modes disappear, but instead our system hosts states localized at specific locations in the bulk of the sample (see Fig. 8). We attribute their existence to the lack of translational symmetry. Indeed, in periodic crystals, isolated localized states cannot exist within the bulk of the material, but this does not have to be the case in a material which lacks translational symmetry. The same thing from a different perspective: in a periodic crystal, corner modes can be viewed as a consequence of bulk polarization [25], which is uniform in the bulk. However, bulk polarization does not have to be uniform in a system without translational symmetry. We leave the problem of building a more comprehensive theory of localized states in quasicrystals for future work.
@@
Figure 2: Schematic phase diagram of the $30^{\circ}$ degrees twisted bilayer Haldane model defined by the Eq. 1 obtained by exact diagonalization. At small $t_{inter}$, the model has exactly the same topological/non-topological phases as a single layer Haldane model. At larger $t_{inter}$, the bulk gap closes, and the in-gap edge states disappear. Simultaneously, another bulk gap may open.
@@
Figure 3: (3(a)) Energies of states plotted against $m$ for various values of $t_{inter}$. We choose $L=30$, $q=2$. The color represents their fractal dimensions. At weak interlayer coupling, the spectrum is gapped for $m>1$, whereas for $m<1$ the gap is filled by the edge states, which have smaller fractal dimensions. At strong inerlayer coupling, the edge states disappear, and the spectrum becomes gapless.
@@
Figure 4: Energies of states plotted against $t_{inter}$ for $m=0$ and $L_{x}=20$ (4(a)), $L_{x}=30$ (4(a)), $L_{x}=30$ (4(c)) as well as $m=0.5$ (4(a)), $m=1.0$ (4(a)), $m=1.5$ (4(c)) at $L_{x}=30$. One can see that at small $t_{inter}$, the system retains properties of its counterpart at $t_{inter}=0$, but at large $t_{inter}$, the system either become gapless ($m=0.5,1$), or a new gap opens ($m=0,1.5$). Localized states have the smallest fractal dimensions and thus are shown in blue color.
@@
Figure 5: Energies (vertical axis) and fractal dimensions (color) of the states plotted against the system size $L_{x}$ for $t_{inter}=0.0$ (5(a)), $t_{inter}=1.0$ (5(b)), $t_{inter}=5.0$ (5(c)) and $t_{inter}=10.0$ (5(d)). States with fractal dimensions smaller than $0.75$ are emphasized by larger dot size. One can see that at $t_{inter}=0.0$ and $t_{inter}=1.0$, topological edge states, as well as localized states within the bulk band exist for any $L_{x}$. In the strongly coupled phase ($t_{inter}=5,10$), one can see a new energy gap, but with increasing $L_{x}$ it becomes less pronounced. The most localized states (blue) may or may not lie within the energy gap.
@@
Figure 6: Energies of states plotted against their fractal dimensions for $m=0$, $L_{x}=30$, $t_{inter}=0$ (6(a)), 11 (6(b)), 55 (6(c)) and $L_{x}=40$, $t_{inter}=0$ (6(d)), 11 (6(e)), 55 (6(f)). One can see that at $t_{inter}=0$ (6(a), 6(d)), the system forms a Chern insulator with topological edge states in the bulk gap, but in addition possesses edge states within the bulk band. As $t_{inter}$ increases (6(b), 6(e)), some of the edge states become localized near the corners. Such corner states persist at strong interlayer coupling (6(c), 6(f)). In addition, states localized at the center of the lattice emerge.
@@
Figure 7: $|\psi_{i}|^{2}$ for a few of the localized states at $m=0$, $t_{inter}=5$, $L_{x}=30$. We observed states localized at the center of the lattice (7(a), 7(b)), forming a ring (7(c), 7(d)), a well as localized at the corners (7(e)) and at the edges (7(f))
@@
Figure 8: $|\psi_{i}|^{2}$ for a few of the most localized localized states at $m=1.5$, $t_{inter}=5$, $L_{x}=30$. One can see that the states can be localized at various locations of the lattice.

### v4

@@ Abstract
We consider $30^{\circ}$ twisted bilayer formed by two copies of Haldane model and explore its evolution with varying interlayer coupling strength. Specifically, we compute the system’s energy spectrum, its fractal dimensions, topological entanglement entropy, local Chern markers and anomalous Hall conductivity. We find that at weak interlayer coupling, the system still has a bulk energy gap, topological edge states and retains topological properties of the isolated layers, but at strong interlayer coupling, this energy gap closes. However, at small values of the Haldane mass $m$, another bulk gap opens. At strong interlayer coupling, the system possesses multiple states localized at various locations of the lattice, including corner states. We emphasize that these corner states do not originate from the topological edge states at the weak coupling, and their location is not necesarily attributed to the bulk gap. We also compute fractal dimensions and establish that the system at large interlayer coupling is multifractal. Finally, we establish that topological entanglement entropy and anomalous Hall conductivity can be used to characterize the system’s topological properties in the same way as a local Chern marker. Our results suggest that the bulk gap at the strong interlayer coupling has non-topological origin.
@@ I Introduction
Motivated by this, we are interested in studying a quasicrystal obtained by smoothly deforming a crystal with known non-trivial topology. Specifically, we choose to consider 30 degrees twisted bilayer honeycomb lattice - a model which is currently actively studied in the context of twisted bilayer graphene [17, 26, 16, 27]. However we consider a twisted bilayer, in which each monolayer is described by gapped Haldane model [21, 23, 24]. We explore our model’s evolution once the interlayer coupling is turned on. We obtain that at weak coupling, the model retains all properties of Haldane model: its bulk spectrum remains gapped and the topological edge states persist. Thus we show that our system of two coupled copies of Chern insulators can be viewed as a topological quasicrystal. On the other hand, as the interlayer coupling becomes larger, the bulk energy gap closes, and the topological edge states disappear. Yet, since this system is non-periodic, its eigenstates are generally not extended. We obtain numerically that our bilayer possesses multiple kinds of states localized at various locations of the lattice. These include, but not limited to corner states, which were previously found in Refs. [23, 24]. We also observe that at large interlayer coupling, a new energy gap may open, but we argue that it is not topological.
@@
This paper is organized as follows. In Sec. II we introduce our model, and in Sec. II.1 we describe its phase diagram, which consists of multiple phases: topological weakly-coupled, non-topological weakly coupled and strongly coupled multifractal. The latter may have both gapped and gapless phases. We also comment on multifractal properties of our model (Sec. II.2). In Sec. II.3 we describe its topological properties. Specifically, in Sec. II.3.1 we present our results for its topological entanglement entropy and in Sec. II.3.2 we describe its topological properties using the ’local Chern markers’. In Sec. II.3.3 we discuss anomalous Hall conductivity. We summarize our findings in Sec. III .
@@ II Model
$H=\sum\limits_{\alpha=1,2}H_{\alpha}+V_{1-2},$ (1)
where $\alpha$ is the layer index, $H_{\alpha}$ is Haldane Hamiltonian of each layer, and $V_{1-2}$ describes interaction between the layers. Specifically,
@@
Here $i,j$ numerate the lattice sites within each layer, <i,j><i,j> is a sum over the nearest neighbors, $\ll i,j\gg$ is a sum over the next-nearest neighbors. $t_{intra}$ is the nearest neghbors hopping (between A, B sites), $t_{2}$ is the next-nearest neighbors hopping (between A-A or B-B sites), $\nu_{ij}=\pm 1$ depending on the relative direction between $i$ and $j$ (see Fig. 2), $\phi=\pi/2$, $m$ is Haldane mass, which we assume to be non-negative, and $\epsilon_{i}=\pm 1$ for A(B) sites (see e.g. ref. [4] for more details). The interlayer potential has the form
$v_{i,j}=\left\{e^{-\frac{|\vec{r}_{i,1}-\vec{r}_{j,2}|}{r_{0}}},\quad|\vec{r}_{i,1}-\vec{r}_{j,2}|<r_{max}\atop 0,\qquad\qquad\quad|\vec{r}_{i,1}-\vec{r}_{j,2}|>r_{max}.\right$. (4)
Here $\vec{r}_{i,1},\vec{r}_{j,2}$ denote the positions of the lattice sites $i,j$ located at the layers $1,2$, the distance between them is $|\vec{r}_{i,1}-\vec{r}_{j,2}|$, and to simplify the model we neglect interlayer spacing. Also we introduced long-distance cutoff $r_{max}$ to decrease the complexity of our numerical calculations. We fix the numerical values of our parameters as $t_{intra}=1$, $t_{2}=1$, $\phi=\pi/2$, $r_{0}=1$, $r_{max}=2$ (we checked that changing $t_{2}$, $r_{max}$ does not have qualitative impact on the model’s properties) and explore its evolution over $m$, $t_{inter}$. Since our 30 degrees twisted bilayer is a non-periodic system, we consider a system forming a square with equal lengths in both directions $L_{x}=L_{y}$ and fix open boundary conditions.
@@
Figure 1: A schematic picture of the $30^{\circ}$ twisted bilayer of two honeycomb lattices. Here we show its top-left corners with $L_{x,y}=20$ (1(a)), $L_{x,y}=30$ (1(b)), $L_{x,y}=40$ (1(c)). The top layer (blue) is formed by sites with coordinates $\vec{r}_{A,B}=\vec{a}_{1}n_{1}+\vec{a}_{2}n_{2}+\vec{c}_{A,B}$, where $\vec{a}_{1}=(\sqrt{3},0)$ , $\vec{a}_{2}=(\sqrt{3}/2,3/2)$, $\vec{c}_{A,B}=(0,\pm 1)$. The bottom layer (red) is obtained by rotating the top layer at 30 degrees. We consider a system with open boundary conditions, whose sites are located inside a square $L_{x,y}<x,y<L_{x,y}$. The cases (1(a), 1(b)) correspond to zigzag-armchair edges, and the case (1(c)) corresponds to the bearded-armchair edge.
((a))
((b))
Figure 2: A schematic picture of couplings in Haldane model in the top (2(a)) and bottom (2(b)) layers. In each layer, there is a nearest neighbors coupling $t$ between AA, $B$ sites and a next nearest neighbors coupling $t_{2}e^{\pm i\phi}$. The arrows mark the sign of each coupling, namely an arrow pointing from the site $i$ to the site $j$ means that the corresponding term in the Hamiltonian has the form $t_{2}e^{+i\phi}c_{i}^{\dagger}c_{j}$. The bottom layer can be obtained by rotating the top layer by $30^{\circ}$ degrees.
@@ II.1 Phase diagram
where $P_{q}$ is a participation ratio defined as $P_{q}=\sum_{i}(\psi^{\dagger}_{i}\psi_{i})^{q}$, and $i=1\ldots N$ runs over all lattice sites. Here we consider a few fixed values of $q$, but unless specified otherwise, we assume $q=2$. In a system of large spatial size, fractal dimension is expected to be close to $1$ for extended bulk states, $1/2$ for edge states and 0 for localized states (e.g. corner modes). In this section, we use $D_{q}$ to find the most localized states, which have the smallest $D_{q}$.
@@
First, let us recall our model’s properties in the known case when the interlayer coupling $t_{inter}$ is set to zero. This model is a Chern insulator at $m<1$ (we assume that both layers have the same signs of Chern numbers). Its bulk states form two bands separated by an energy gap, which in turn, is filled by edge states. At $m=1$, the bulk gap closes, and at $m>1$, it reopens again, and the system becomes a trivial insulator.
@@
Once the interlayer coupling is turned on, the phase diagram of the system in $m,t_{inter}$ space has multiple phases, which we schematically outline on the Fig. 3 and which we obtained by running over different values of $m$ and $t_{inter}$ (see Figs. 4, 5). Specifically, if we start from the topological phase of the Haldane model, we observe (see Figs. 4) that at small coupling, the bulk states are still separated by an energy gap and the edge states persist, but as $t_{inter}$ is increasing, the bulk energy gap shrinks. Once $t_{inter}$ reaches a critical value ($t_{inter}\approx 1.5$ for $m=0$), the bulk gap closes, and the edge states disappear. As we increase $t_{inter}$ further, at small $m$, another energy gap appears (see Fig. 5(a), 5(b), 5(c) ). However, we do not observe it at larger $m$ (see Fig. 5(d) ). Similarly, if we start from the non-topological phase (see Fig. 5(f)), at small $t_{inter}$, the system remains gapped, and as $t_{inter}$ is increasing, the gap becomes smaller and eventually disappears. Near this phase transition, another gap may appear.
@@
Let us look more closely at the properties of the states in our model. For simplicity, we focus on the case $m=0$. As we mentioned, at $t_{inter}=0$, our model is a superposition of two Haldane models, which in turn host bulk states separated by an energy gap and edge states within the bulk gap. However, we also found that our system hosts additional edge states within the bulk gap. We observe such states in the cases of both zigzag-armchair and bearded-armchair edges, but we find them to lie at slightly different energies. Specifically, in the case of zigzag-armchair edges, these edge states lie at $E\approx\pm 2.3$ (see Fig. 7(a)), whereas in the case of bearded-armchair edge, these states lie at energy $E\approx\pm 1.9$. One can see these states for various system sizes $L_{x,y}$ on the Fig. 6(a) - the bearded-armchair configuration corresponds to lattice sizes $7,10,13,16\ldots$, and zigzag-armchair is realized for all other integer values of $L_{x,y}$.
@@
As we increase $t_{inter}$, the edge states in the bottom ‘band’ disappear. However, in the top ‘band’, they persist, and a few of them tend to become localized within the corners of our system (see Figs. 7(b), 7(e)). Specifically for $L_{x,y}=30$, we also find that a few of the topological edge states become localized at the corners (see Fig. 7(b); note that we could not establish connection between these modes and a shape of the lattice boundary). As we increase $t_{inter}$ further, the bulk gap closes, and the system becomes gapless. At even larger $t_{inter}$, another energy gap opens. Interestingly, we found that even though such a gap exists at any system size, its precise location and width varies a lot between different lattice sizes (see Figs. 6(c), 6(d)). We attribute it to the fact that since our system is quasiperiodic, the system becomes unique for each lattice size, and therefore it is extremely difficult to achieve ‘continuum limit’ (see also discussion in the Sec. II.2). We however, can see that as the system size increases, this gap becomes less pronounced (see Fig. 6).
@@
As predicted in the previous works [21], our system hosts states localized at the corners. However, we argue that their location is not attributed to the bulk gap. Indeed, we can see that for $L_{x,y}=30$ (see Fig. 5(b)), the localized corner states are in the bulk gap for a range of $t_{inter}$ between $\approx 5$ and $10$, but at smaller or larger $t_{inter}$, they merge with the bulk spectrum. On the other hand, at $L_{x,y}=20$ (see Fig. 5(a)), we find that the corner modes lie entirely within the bulk spectrum away from the gap, but at the same time, within the bulk gap, there are extended states. We note that both $L_{x,y}=20,30$ in our example correspond to zigzag-armchair edges of the lattice, and even have a similar structure of the corners (see Fig. 1(a), 1(b) ). We observe qualitatively similar picture for $L_{x,y}=40$ with ‘armchair-bearded’ edge structure (Fig. 1(c)): localized corner modes lie within the bulk gap at $t_{inter}\gtrsim 5$, but within the bulk states at smaller $t_{inter}$. ((see Fig. 5(c))).
@@
Even more interestingly, we observe numerically that our system hosts other kinds of localized states. For instance, even at large $t_{inter}$, our system still hosts states localized near edges (see Fig. 8(f) ). But most surprisingly, we find that, at $m=0$, the states with the smallest fractal dimensions are localized in the center of the lattice (see Fig. 8(a), 8(b)).
@@
We also find that at $m\neq 0$, both corner and ‘center’ modes disappear, but instead our system hosts states localized at other specific locations in the bulk of the sample (see Fig. 9). We attribute their existence to the lack of translational symmetry. Indeed, in periodic crystals, isolated localized states cannot exist within the bulk of the material, but this does not have to be the case in a material which lacks translational symmetry. The same thing from a different perspective: in a periodic crystal, corner modes can be viewed as a consequence of bulk polarization [3], which is uniform in the bulk. However, bulk polarization does not have to be uniform in a system without translational symmetry. We leave the problem of building a more comprehensive theory of localized states in quasicrystals for future work.
@@
Figure 3: Schematic phase diagram of the $30^{\circ}$ degrees twisted bilayer Haldane model defined by the Eq. 1 obtained by exact diagonalization. At small $t_{inter}$, the model has exactly the same topological/non-topological phases as a single layer Haldane model. At larger $t_{inter}$, the bulk gap closes, and the in-gap edge states disappear. Simultaneously, another bulk gap opens at $m\approx 0$ and $m\gtrsim 1$.
@@
Figure 4: (4(a)) Energies of states plotted against $m$ for various values of $t_{inter}$. We choose $L_{x}=L_{y}=30$. The color represents their fractal dimensions $D_{q}$ for $q=2$. At weak interlayer coupling, the spectrum is gapped for $m>1$, whereas for $m<1$ the gap is filled by the edge states, which have smaller fractal dimensions. At strong inerlayer coupling, the edge states disappear, and the spectrum becomes gapless.
@@
Figure 5: Energies of states plotted against $t_{inter}$ for $m=0$ and $L_{x,y}=20$ (5(a)), $L_{x,y}=30$ (5(b)), $L_{x,y}=40$ (5(c)) as well as $m=0.5$ (5(a)), $m=1.0$ (5(a)), $m=1.5$ (5(c)) at $L_{x,y}=30$. One can see that at small $t_{inter}$, the system retains properties of its counterpart at $t_{inter}=0$, but at large $t_{inter}$, the system either become gapless ($m=0.5,1$), or a new gap opens ($m=0,1.5$). The color represents each state’s fractal dimension $D_{q}$ at $q=2$, and we use the same colorbar as on the Fig. 4. Localized states have the smallest fractal dimensions and thus are shown in blue color, whereas more extended states are shown in green.
@@
Figure 6: Energies (vertical axis) and fractal dimensions (color) of the states plotted against the system size $L_{x}=L_{y}$ for $t_{inter}=0.0$ (6(a)), $t_{inter}=1.0$ (6(b)), $t_{inter}=5.0$ (6(c)) and $t_{inter}=10.0$ (6(d)). The color represents each state’s fractal dimension $D_{q}$ at $q=2$, and we use the same colorbar as on the Fig. 4. States with fractal dimensions smaller than $0.75$ are emphasized by larger dot size. One can see that at $t_{inter}=0.0$ and $t_{inter}=1.0$, topological edge states, as well as localized states within the bulk band exist for any $L_{x}$. In the strongly coupled phase ($t_{inter}=5,10$), one can see a new energy gap, but with increasing $L_{x,y}$ it becomes less pronounced. The most localized states (blue) may or may not lie within the energy gap.
@@
Figure 7: Energies of states plotted against their fractal dimensions for $m=0$, $L_{x,y}=30$, $t_{inter}=0$ (7(a)), $t_{inter}=1$ (7(b)), $t_{inter}=5$ (7(c)) and $L_{x,y}=40$, $t_{inter}=0$ (7(d)), $t_{inter}=1$ (7(e)), $t_{inter}=5$ (7(f)). One can see that at $t_{inter}=0$ (7(a), 7(d)), the system forms a Chern insulator with topological edge states in the bulk gap, but in addition possesses edge states within the bulk band. As $t_{inter}$ increases (7(b), 7(e)), some of the edge states become localized near the corners. Such corner states persist at strong interlayer coupling (7(c), 7(f)). In addition, states localized at the center of the lattice emerge.
@@
Figure 8: $|\psi_{i}|^{2}$ for a few of the localized states at $m=0$, $t_{inter}=5$, $L_{x,y}=30$. We observed states localized at the center of the lattice (8(a), 8(b)), forming a ring (8(c), 8(d)), a well as localized at the corners (8(e)) and at the edges (8(f))
@@
Figure 9: $|\psi_{i}|^{2}$ for a few of the most localized localized states at $m=1.5$, $t_{inter}=5$, $L_{x,y}=30$. One can see that the states can be localized at various locations of the lattice.